# Beyond Memorization: The Challenge of Random Memory Access in Language Models

### v2

@@ Abstract
[…]
@@ 1 Introduction
Language models (LMs) have recently showcased outstanding abilities in NLP tasks with a large amount of memory stored in their parameters (Brown et al., 2020; Ouyang et al., 2022). Through pre-training on large text corpora, LMs memorize factual knowledge about the world (Zhou et al., 2023). Consequently, they show great performance in knowledge-intensive tasks (Petroni et al., 2021) such as open-domain question answering (Kamalloo et al., 2023; Ziems et al., 2023; Mallen et al., 2023). There is a growing interest in considering LMs as knowledge bases (Wang et al., 2021; Heinzerling and Inui, 2021; Petroni et al., 2019; Cao et al., 2021; AlKhamissi et al., 2022). Despite the recent advances in applying LMs to solve downstream tasks, the fundamentals of how LMs store knowledge and access memory in their parameters remains a subject of ongoing research and intrigue (Tirumala et al., 2022; Zhu and Li, 2023; Allen-Zhu and Li, 2023; Berglund et al., 2023).
Figure 1: A illustration of our investigation of memory access pattern in language models. We find that the model accesses its parametric memory largely in a sequential manner, and faces difficulty in randomly accessing the content in the middle of memorized strings.
@@
With these concepts, we design experiments with both synthetic and real data to evaluate the language model’s ability to perform sequential or random access to memorized content, as illustrated in Figure 1. We limit our study to decoder-only language models due to their increasing popularity and capability (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a, b; Jiang et al., 2023). We first ask the model to memorize key–value pairs of various types and show that the model is able to sequentially read memorized content to a satisfying degree. Next, we test the model’s random access ability by training it to recite a sentence or find an answer to a question in a memorized passage. In such tasks, the model’s performance falls drastically when it is required to extract a span in the middle of a passage, revealing its incapability to randomly access its memory.
@@
  • We show that language models can access their memory sequentially and can reproduce memorized content. In contrast, they encounter significant challenges in random memory access.
@@ 2 Related Work
Previous works have shown that using prompts can effectively retrieve knowledge stored in large language models (Bouraoui et al., 2019; Jiang et al., 2021; Wang et al., 2021). We follow earlier work to use prompts to query the model to access and regenerate memorized content. However, a notable difference is that prior work focuses on finding optimised methods to elicit the model’s knowledge obtained during pretraining (Youssef et al., 2023; Liu et al., 2023), while we directly use unique keys for memorizing and retrieving content.
@@ 3.1 Task Formulation
We abstract the language model as a memory bank and investigate its sequential or random access ability. We adopt a simple definition of a memory bank as a key–value store $\mathcal{D}=\{k_{i}:p_{i}\}$, where $k_{i}$ represents a unique identifier (ID) assigned to the content of the $i$-th passage [^1] We use the term “document” and “passage” interchangeably in this paper, referring to a chunk of text..
There are two core functions that a memory bank needs to support: reading and writing.
Given that our memory bank is embodied as a language model, it is not straightforward to write and read the model’s memory. Following previous work (Zhu and Li, 2023; Wang et al., 2021), for writing to the memory bank, we use fine-tuning to update the model’s parameters. For reading, we use prompting to elicit the model’s memory. Specifically, for each passage $p_{i}$ with its corresponding identifier $k_{i}$, we create two types of data instances: writing, $S_{write}(k_{i},p_{i})$ and reading, $S_{read}(k_{i})\rightarrow p_{i}$, where $S_{write}$ and $S_{read}$ denote the prompts detailed in Appendix A.1.
As the primary goal of our study is to test whether the model can read (access) its stored content sequentially or randomly, we mainly vary the reading function across different experiments.
Given a corpus consisting of $M$ passages, we split the corpus into two subsets: $T$ training passages and $V=M-T$ validation passages. We adopt a mixed training strategy as described by Zhu and Li (2023): During the training stage, we include $S_{write}$ and $S_{read}$ instances of T training passages, as well as $S_{write}$ instances of V validation passages. Our objective is for the model to learn to associate each identifier with its passage content by training on the reading and writing instances of the training passages. During evaluation, we prompt the model with the $S_{read}$ instances of the V validation passages to test the model’s memory access pattern.
@@ Discussion.
Figure 2: EM and BLEU for reading validation passages, varying the number of training passages. We calculate EM using only the first 25 tokens, as the model often generates beyond the max passage length (25).
@@
We observe that even with a training passage count of 50k, the model could accurately reproduce over 70% of memorized validation passages. However, there is also a bottleneck in parametric memory: the performance drops to nearly zero when the passage count exceeds 100k. We attribute this bottleneck to the difficulty in training, as the model fails to converge on memorizing all the passages. Therefore, in subsequent experiments, we carefully manage the corpus size to ensure that the model memorizes all passages.
@@ 3.3 Random Access: Selective Recitation
We follow Mallick et al. (2023) to place markers at the boundaries of each sentence, obtained by the NLTK sentence splitter [^3] https://www.nltk.org/api/nltk.tokenize.sent_tokenize.html: a passage is formatted as “[0] sent0 [0] [1] sent1 [1], …,”. In this case, the model only needs to learn to copy the content between these markers. Our selective recitation task requires the model to recite the $j$-th sentence of passage $p_{i}$ based on the given passage ID $k_{i}$. The reading function is now $S_{read}(k_{i},j)\rightarrow p_{i}[j]$, such as “What is sentence [1] of Document #2033?”. For reference, we also test the model’s performance in a baseline where the passage content is provided in the context window.
@@
Figure 3: A stacked bar plot showing the accuracy of ID-guided sentence recitation with different marker numbers. The performance decreases significantly as the sentence index grows.
@@
w.o. passage memorization
Closed-Book QA 9.0 16.6 9.0 16.6 9.0 16.6
Open-Book QA 73.7 79.3 73.7 79.3 73.7 79.3
w. passage memorization
Grounded QA w. Golden ID 26.7 35.6 20.7 28.7 24.3 32.6
@@ 3.4 Random Access: Grounded Question Answering
We experiment with the well-known SQuAD-v1 (Rajpurkar et al., 2016) dataset because many of its questions are closely dependent on the passage, such as “How did the war start?”. This design compels the model to depend on the memorized IDs and passages rather than pre-existing knowledge. We explore the grounded QA task with variants of providing (1) the ID of the golden passage with the answer, (2) a random non-golden ID and (3) no ID. For comparison, we also consider the setups that do not involve writing passages to the model’s parametric memory. These include (1) closed-book QA, where the model is fine-tuned solely on QA pairs, serving as a baseline to assess the model’s reliance on prior knowledge for answering questions, and (2) open-book QA, where the golden passage content is concatenated with the question, setting the upper limit of extractive QA performance.
@@ 4 Mitigating Random Access Challenge
The earlier experiments show that in general, language models perform well in sequentially accessing their parametric memory, but encounter challenges in random memory access. This naturally raises the question: How can we mitigate the shortcomings in random memory access?
@@ 4.1 Proposed Method
First, we add a setup to the selective sentence recitation task: Based on the given ID, the model is tasked to first recite the entire content of the corresponding passage and then the specific sentence, altering the reading operation to $S_{read}(k_{i},j)\rightarrow(p_{i},p_{i}[j])$. Similarly, for the grounded QA task, we ask the model to recite the passage associated with the input passage ID before answering the question. In the setup without an ID, the model is still trained to recite the golden passage.
To explore the effect of permutation during the writing stage, we perform permutation among sentences in a passage to create diverse $S_{write}$ instances. For a $J$-sentence passage we tested: (1) first, moving each sentence to the passage’s beginning to create $J$ unique instances;
(2) random, randomly shuffling the sentences $k$ times to create $k$ instances, where $k$ is set to 4 by default;
@@ Discussion.
Similarly, explicitly reciting the golden passages markedly enhances question-answering performance, as shown in Table 2. This observation is consistent across all three types of passage IDs. Conversely, intentionally prompting the model to recite a random passage leads to a decline in performance. This is likely because random passages introduce irrelevant information and confuse the model. Surprisingly, the recitation of relevant passages benefits performance even without an ID, although the improvement is smaller than with the golden ID. This verifies the effectiveness of recitation in more general settings of question answering.
Another way of enhancing random access is to perform permutation of sentences, as presented in Table 3. Simply bringing every sentence to the start of the passage once (first) or randomly permuting the sentences many times (random) helps to solve the challenge of accessing the middle content of a passage. We also find that permutation enhances grounded QA performance (Appendix C) from 26.7 to 31.3 in terms of EM. However, it is noteworthy that permutation does not alter the inherent sequential access pattern of parametric memory. Rather, by permuting the sentences and disrupting their original order, we allow more sentences to be sequentially accessible via the ID.
@@ 5 Case Study: Open-Domain Question Answering
Closed-Book QA w. Mix Training 12.6 18.2 - 15.7 22.8
@@ 5.1 Experimental Setup
We start from a baseline setup where the training only involves QA pairs, i.e., closed-book QA. Next, we consider two types of training strategies to write the passages into the memory. In the mixed setting, the model is fine-tuned on a mixture of the $S_{write}$ instances of all passages and training QA pairs. In the continual setting, the model is fine-tuned on $S_{write}$ of all passages first, followed by fine-tuning on QA. To test the effectiveness of recitation, we also include settings where the model is trained to recite the golden passage(s) before answering.
As the task requires the model to perform both passage retrieval and question answering, we expect that the model size should be sufficiently large. Therefore, we opt for GPT2-XL with 1.5B parameters. In the mixed setting, we train the model for 20 epochs with a learning rate of 3e-5. In the continual setting, we first train 20 epochs on the passages, followed by another 20 epochs on QA pairs. We report the best performance based on the EM score on validation questions.
@@ 5.2 Results and Discussion
The results presented in Table 4 demonstrate that writing golden passages into the model’s memory, with either mixed or continual training, leads to improved performance over the baseline closed-book setting. This aligns with our expectations, as we deliberately inject passages containing the answers to the questions into the memory, enriching the model’s knowledge.
@@
The mixed training strategy outperforms the continual training setup. This is likely because the model’s memory of passage content is refreshed constantly in mixed training. In continual training, however, the second stage only involves QA pairs on training passages, potentially leading to fading memory of validation passages. Consequently, the recitation becomes less accurate, as shown by a decrease in the BLEU score.
@@ 6 Conclusion
In this work, we mainly explore the memory access pattern of decoder-only language models of the GPT2 family. Future research is needed to understand whether our conclusions apply to other types of language models based on transformers such as encoder-only models and encoder-decoder models. Furthermore, we do not extend our study to larger models beyond 1.5B parameters due to computing resource constraints. It might be worthwhile to explore the scaling behavior of memory access patterns in larger language models. In addition, we mainly conduct controlled experiments on a text corpus of fixed size. Further investigation may be needed to explore how the findings can apply to large-scale pretraining corpus and their implications on pretrained language models.
@@ Acknowledgements
We appreciate the suggestions and comments by Do Xuan Long, Yanxia Qin, Yisong Miao, and other members in NUS WING. Tongyao Zhu is supported by the Industry PhD Program of Sea AI Lab.
@@ B.1 Reciting the first/second/last sentence
Table 5: BLEU and EM score of reciting the first, second or last sentence of a memorized passage.
@@ B.2 Reciting the next/previous sentence
Table 6: BLEU and EM score of reciting the next or previous sentence given an input sentence.
@@
We observe that finding the sentence after the input sentence is always easy, while the reverse task is much more difficult. This also reveals that the model reads its memory sequentially. It is unable to randomly access the sentence before the input $s_{j}$, even if the target sentence is adjacent to the input sentence.
Appendix C Additional Grounded Question Answering Experiments with Permutation
We conduct additional experiments showing the effect of performing permutations of sentences of memorized passages in the grounded question-answering task.
In Table 7, we show the effect of augmenting the $S_{write}$ instances to include sentence permutations of the original passage. We observe that the model’s performance generally improves as we perform permutation. This validates our findings that performing permutation enhances random access to the passage content.
ID=Title ID=Rare ID=Num
Setup EM F1 EM F1 EM F1
Grounded QA w. Golden ID 26.7 35.6 20.7 28.7 24.3 32.6
    $\hookrightarrow$+↪ + Permutation (rand-1) 25.7 35.0 19.0 27.5 19.7 28.1
    $\hookrightarrow$+↪ + Permutation (rand-2) 26.0 35.6 25.7 33.7 23.7 32.8
    $\hookrightarrow$+↪ + Permutation (rand-4) 29.7 38.5 25.3 35.6 25.0 34.2
    $\hookrightarrow$+↪ + Permutation (rand-8) 31.3 40.1 27.7 36.7 29.0 38.3
Table 7: The EM and F1 score of performing sentence permutation during the writing phase. rand-k means that permutation is performed $k$ times.
Appendix D Additional Open-Domain Question Answering Experiments

### v3

@@ Abstract
Tongyao Zhu1,2 Qian Liu1thanks: Corresponding authors. Liang Pang3∗ Zhengbao Jiang4 Min-Yen Kan2 Min Lin1 1Sea AI Lab  2National University of Singapore 3Institute of Computing Technology, CAS 4Carnegie Mellon University tongyao.zhu@u.nus.edu{liuqian,linmin}@sea.com pangliang@ict.ac.cnzhengbaj@cs.cmu.eduknmnyn@nus.edu.sg
@@ 1 Introduction
Language models (LMs) have recently showcased outstanding abilities in NLP tasks with a large amount of memory stored in their parameters (Brown et al., 2020; Ouyang et al., 2022). Through pre-training on large text corpora, LMs memorize factual knowledge about the world (Zhou et al., 2023). Consequently, they show great performance in knowledge-intensive tasks (Petroni et al., 2021) such as open-domain question answering (Kamalloo et al., 2023; Ziems et al., 2023; Mallen et al., 2023). There is a growing interest in considering LMs as knowledge bases (Wang et al., 2021; Heinzerling and Inui, 2021; Petroni et al., 2019; Cao et al., 2021; AlKhamissi et al., 2022). Despite the recent advances in applying LMs to solve downstream tasks, the fundamentals of how LMs store knowledge and access memory in their parameters remain a subject of ongoing research and intrigue (Tirumala et al., 2022; Zhu and Li, 2023; Allen-Zhu and Li, 2023; Berglund et al., 2023).
Figure 1: An illustration of our investigation of memory access pattern in language models. We find that the model accesses its parametric memory largely in a sequential manner, and faces difficulty in randomly accessing the content in the middle of memorized strings.
@@
With these concepts, we design experiments with both synthetic and real data to evaluate the language model’s ability to perform sequential or random access to memorized content, as illustrated in Figure 1. We limit our study to decoder-only language models because of their increasing popularity and capability (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a, b; Jiang et al., 2023). We first ask the model to memorize key–value pairs of various types and show that the model is able to sequentially read memorized content to a satisfying degree. Next, we test the model’s random access ability by training it to recite a sentence or find an answer to a question in a memorized passage. In such tasks, the model’s performance falls drastically when it is required to extract a span in the middle of a passage, revealing its incapacity to randomly access its memory.
@@
  • We show that language models can access their memory sequentially and can reproduce memorized content, but encounter significant challenges in random memory access.
@@ 2 Related Work
Previous works have shown that using prompts can effectively retrieve knowledge stored in large language models (Bouraoui et al., 2019; Jiang et al., 2021; Wang et al., 2021). We follow earlier work to use prompts to query the model to access and regenerate memorized content. However, a notable difference is that prior work focuses on finding optimised methods to elicit the model’s knowledge obtained during pretraining (Youssef et al., 2023; Liu et al., 2023; Yu et al., 2023), while we directly use unique keys for memorizing and retrieving content.
@@ 3.1 Task Formulation
We abstract the language model as a memory bank and investigate its sequential or random access ability. We adopt a simple definition of a memory bank as a key–value store $\mathcal{D}=\{k_{i}:p_{i}\}$, where $k_{i}$ represents a unique identifier (ID) assigned to the content of the $i$-th passage [^1] We use “document” and “passage” interchangeably to refer to a chunk of text..
There are two core functions that a memory bank needs to support: reading and writing. Given that our memory bank is embodied as a language model, it is not straightforward to write and read the model’s memory. Following previous work (Zhu and Li, 2023; Wang et al., 2021), for writing to the memory bank, we use fine-tuning to update the model’s parameters. For reading, we use prompting to elicit the model’s memory. Specifically, for each passage $p_{i}$ with its corresponding identifier $k_{i}$, we create two types of data instances: writing, $S_{write}(k_{i},p_{i})$ and reading, $S_{read}(k_{i})\rightarrow p_{i}$, where $S_{write}$ and $S_{read}$ denote the prompts detailed in Appendix A.1.
As the primary goal of our study is to test whether the model can read (access) its stored content sequentially or randomly, we vary the reading function across different experiments. Given a corpus consisting of $M$ passages, we split the corpus into two subsets: $T$ training passages and $V=M-T$ validation passages. We adopt a mixed training strategy as described by Zhu and Li (2023): During the training stage, we include $S_{write}$ and $S_{read}$ instances of T training passages, as well as $S_{write}$ instances of V validation passages. Our objective is for the model to learn to associate each identifier with its passage content by training on the reading and writing instances of the training passages. During evaluation, we prompt the model with the $S_{read}$ instances of the V validation passages to test the model’s memory access pattern.
@@ Discussion.
Figure 2: EM and BLEU for reading validation passages, varying the number of training passages. We calculate EM using only the first 25 tokens, as the model tends to continue generation beyond the max passage length (25).
@@
We observe that even with a training passage count of 50k, the GPT2-large model accurately reproduces over 70% of memorized validation passages. However, there is also a bottleneck in parametric memory: the performance drops to nearly zero when the passage count exceeds 100k. We attribute this bottleneck to the difficulty in training, as the model fails to converge on memorizing all the passages. Therefore, in subsequent experiments, we carefully manage the corpus size to ensure that the model memorizes all passages.
@@ 3.3 Random Access: Selective Recitation
We follow Mallick et al. (2023) to place markers at the boundaries of each sentence, obtained by the NLTK sentence splitter [^3] https://www.nltk.org/api/nltk.tokenize.sent_tokenize.html: a passage is formatted as “[0] sent0 [0] [1] sent1 [1], …,”. In this case, the model only needs to learn to copy the content between these markers. Our selective recitation task requires the model to recite the $j$-th sentence of passage $p_{i}$ based on the given passage ID $k_{i}$. The reading function is now $S_{read}(k_{i},j)\rightarrow p_{i}[j]$, such as “What is sentence [1] of Document #2033?” shown in Figure 1. For reference, we also test the model’s performance in a baseline where the passage content is provided in the context window.
@@
Figure 3: A stacked bar plot showing the accuracy of ID-guided sentence recitation with different marker numbers. The performance decreases significantly as the sentence index grows, revealing the model’s incapability in accessing middle sentences.
@@
w/o passage memorization
Closed-Book QA (lower bound) 9.0 16.6 9.0 16.6 9.0 16.6
Open-Book QA (upper bound) 73.7 79.3 73.7 79.3 73.7 79.3
w/ passage memorization
Grounded QA w/ Golden ID 26.7 35.6 20.7 28.7 24.3 32.6
@@ 3.4 Random Access: Grounded Question Answering
We experiment with the well-known SQuAD-v1 (Rajpurkar et al., 2016) dataset because many of its questions are closely dependent on the passage, such as “How did the war start?”. Without reference to an article, the question can be ambiguous and unanswerable. This design compels the model to depend on the memorized IDs and passages rather than pre-existing knowledge. We explore the grounded QA task with variants of providing (1) the ID of the golden passage with the answer, (2) a random non-golden ID and (3) no ID. For comparison, we also consider the setups that do not involve writing passages to the model’s parametric memory. These include (1) closed-book QA, where the model is fine-tuned solely on QA pairs, serving as a lower-bound baseline to assess the model’s reliance on prior knowledge for answering questions, and (2) open-book QA, where the golden passage content is concatenated with the question, setting the upper limit of extractive QA performance.
@@ 4 Mitigating Random Access Challenge
Our earlier experiments show that in general, language models perform well in sequentially accessing their parametric memory, but encounter challenges in random memory access. This naturally raises the question: How can we mitigate the shortcomings in random memory access?
@@ 4.1 Proposed Method
First, we add a setup to the selective sentence recitation task: Based on the given ID, the model is tasked to first recite the entire content of the corresponding passage and then the specific sentence, altering the reading operation to $S_{read}(k_{i},j)\rightarrow(p_{i},p_{i}[j])$. Similarly, for the grounded QA task, we ask the model to recite the passage associated with the input passage ID, before answering the question. In the setup without an ID, the model is still trained to recite the golden passage.
To explore the effect of permutation during the writing stage, we perform permutation among sentences in a passage to create diverse $S_{write}$ instances. For a $J$-sentence passage, we test: (1) first, moving each sentence to the passage’s beginning to create $J$ unique instances; (2) random-k, randomly shuffling the sentences $k$ times to create $k$ instances, where $k$ is set to 4 by default. To show that the effect of permutation is not simply due to more training data, we also include a baseline dup-J, where each passage is duplicated $J$ times in training data.
@@ Discussion.
Similarly, explicitly reciting the golden passages markedly enhances question-answering performance, as shown in Table 2 (+Recitation). This observation is consistent across all three types of passage IDs. Conversely, intentionally prompting the model to recite a random passage leads to a decline in performance. This is likely because random passages introduce irrelevant information and confuse the model. Surprisingly, the recitation of relevant passages benefits performance even without an ID, although the improvement is smaller than with the golden ID. This verifies the effectiveness of recitation in more general settings of question answering.
Another way of enhancing random access is to perform permutation of sentences, as presented in Table 3. Simply bringing every sentence to the start of the passage once (first) or randomly permuting the sentences many times (random) helps to solve the challenge of accessing the middle content of a passage. In contrast, simply duplicating the original passage does not contribute to enhanced random access. We also observe that permutation during writing time enhances grounded QA performance (Table 4), which monotonically increases with the number of random permutations. However, it is noteworthy that permutation does not alter the inherent sequential access pattern of parametric memory. Rather, by permuting the sentences and disrupting their original order, we allow more sentences in the middle or the end of a passage to be sequentially accessible via the ID.
We also verify that the conclusions are generalizable to larger decoder-only LMs. In Appendix D, we observe similar challenges in random access in Qwen1.5-4b (Bai et al., 2023), and Llama2-7B (Touvron et al., 2023b) across different tasks. Moreover, we observe that such challenges could be effectively mitigated by our proposed methods of recitation and permutation.
Title (ID Type) Rare (ID Type) Num (ID Type)
Setup EM F1 EM F1 EM F1
Grounded QA w. Golden ID 26.7 35.6 20.7 28.7 24.3 32.6
      $\hookrightarrow$+↪ + Duplication (dup-J) 26.7 36.8 20.7 28.3 22.3 30.6
      $\hookrightarrow$+↪ + Permutation (random-1) 25.7 35.0 19.0 27.5 19.7 28.1
      $\hookrightarrow$+↪ + Permutation (random-2) 26.0 35.6 25.7 33.7 23.7 32.8
      $\hookrightarrow$+↪ + Permutation (random-4) 29.7 38.5 25.3 35.6 25.0 34.2
      $\hookrightarrow$+↪ + Permutation (random-8) 31.3 40.1 27.7 36.7 29.0 38.3
Table 4: The EM and F1 score of performing sentence permutation during the writing phase. random-k means that permutation is performed $k$ times.
@@ 5 Case Study: Open-Domain Question Answering
Closed-Book QA w. Mixed Training 12.6 18.2 - 15.7 22.8
@@ 5.1 Experimental Setup
We start from a baseline setup where the training only involves QA pairs, i.e., closed-book QA, which evaluates the model’s prior knowledge gained from pretraining. Next, we consider two types of training strategies to write the passages into the memory. In the mixed setting, the model is fine-tuned on a mixture of the $S_{write}$ instances of all passages and training QA pairs. In the continual setting, the model is fine-tuned on $S_{write}$ instances of all passages first, followed by fine-tuning on QA. To test the effectiveness of recitation, we also include settings where the model is trained to recite the golden passage(s) before answering.
As the task requires the model to perform both passage retrieval and question answering, we expect that the model size should be sufficiently large. Therefore, we upgrade our LLM to GPT2-XL with 1.5B parameters. In the mixed setting, we train the model for 20 epochs with a learning rate of 3e-5. In the continual setting, we first train 20 epochs on the passages, followed by another 20 epochs on QA pairs. We report the best performance based on the EM score on validation questions.
@@ 5.2 Results and Discussion
Table 5 demonstrates that writing golden passages into the model’s memory leads to improved performance over the baseline closed-book setting, with either mixed or continual training. This aligns with our expectations, as we deliberately inject passages containing the answers to the questions into the memory, enriching the model’s knowledge.
@@
The mixed training strategy outperforms the continual training setup. This is likely because the model’s memory of passage content is constantly refreshed in mixed training. In contrast, during continual training, the second stage only involves QA pairs on training passages, potentially leading to fading memory of validation passages. Consequently, the recitation becomes less accurate, as shown by a decrease in the BLEU score.
@@ 6 Conclusion
In this work, we mainly explore the memory access pattern of decoder-only language models. Future research is needed to understand whether our conclusions apply to other types of language models based on transformers such as encoder-only models and encoder-decoder models. Furthermore, we do not extend our study to larger models beyond 7 billion parameters due to computing resource constraints. It might be worthwhile to explore further scaling behavior of memory access patterns in even larger language models. In addition, we mainly conduct controlled experiments on a text corpus of fixed size. Further investigation may be needed to explore how the findings can apply to large-scale pretraining corpus and their implications on pretrained language models.
@@ Acknowledgements
We appreciate the suggestions and comments by Do Xuan Long, Yanxia Qin, Yisong Miao, and other members of NUS WING. Tongyao Zhu is supported by the Industry PhD Program of Sea AI Lab.
@@ B.1 Reciting the first/second/last sentence
Table 6: BLEU and EM scores of reciting the first, second or last sentence of a memorized passage.
@@ B.2 Reciting the next/previous sentence
Table 7: BLEU and EM scores of reciting the next or previous sentence given an input sentence.
@@
We notice that finding the sentence after the input sentence is always easy, while the reverse task is much more difficult. This also reveals that the model reads its memory sequentially. It is unable to randomly access the sentence before the input $s_{j}$, even if the target sentence is adjacent to the input sentence.
Appendix C Additional Open-Domain Question Answering Experiments